How carbon-12 was weighted for the first time? How did they found that 1 mol of carbon 12 is 12g or 6.0221415 * 10^23 atom of carbon is 12g , how did they count? How did they know x atom of carbon is x gram in 19th century






1
Expert's answer
2021-02-22T06:06:27-0500

The carbon-12 atom, which is still used as the standard today, contains six protons and six neutrons for an atomic mass of twelve amu.


The value of the mole is equal to the number of atoms in exactly 12 grams of pure carbon-12. 12.00 g C-12 = 1 mol C-12 atoms = 6.022 × 1023 atoms • The number of particles in 1 mole is called Avogadro's Number (6.0221421 x 1023).



They measured atomic weights in terms of atomic mass units (amu), where 1 amu was equal to one-twelfth of the mass of a carbon-12 atom. When in the second half of the 19th century, chemists used other means to approximate the number of atoms in a given volume of gas — that famous constant known as Avogadro's number — they began producing rough estimates of the mass of a single atom by weighing the volume of the whole gas, and dividing by the number.


And one more thing x atoms of carbon atom is not equal to x gram

1 carbon atom = 12.0107 g

For "x" carbon atom = 12.0107×"x" g